A person on a log flume ride at an amusement park reaches the splashdown area at the end of the ride.  If the mass of the flume log and rider is 190 kg and the flume log has 2800 Joules of kinetic energy as it enters the splashdown, how much work has been done by the water-resistance of the splashdown area to slow down the flume log if it experiences an impulse of -385 N*s?


1
Expert's answer
2021-05-05T19:06:13-0400
"W=K-\frac{p^2}{2m}\\W=2800-\frac{385^2}{2(190)}=2410\ J"